OpenStudy (anonymous):

True or False? Y = 3/x is a Direct variation. Can you explain please

OpenStudy (mertsj):

F. In a direct variation, y is equal to some constant times x not divided by x

OpenStudy (anonymous):

False, this is inverse variation, because in direct (linear) variation, you get something like "y=mx+b," where m is a coefficient, and b is the y-intercept.

OpenStudy (saifoo.khan):

\[y \alpha \frac 1x\]

OpenStudy (anonymous):

Thanks people (:

OpenStudy (mertsj):

yw